# VoiceBridge: Designing Latent Bridge Models for General Speech Restoration at Scale

### v1

@@ 3.4 Perceptual-Aware Generation
After training the EP-VAE encoder $\mathcal{E}$ with Equation 5 for encoding HQ target, its fine-tuned version $\mathcal{E}^{\text{np}}$ with Equation 6 for encoding diverse LQ priors, and the shared decoder $\mathcal{D}$, we can already create a structured latent space where the different LQ priors have converged. Then, we train the bridge transformer model that maps the joint neural prior and the HQ target using Equation 3, while the LQ prior and condition $\mathbf{z}_{1}$ have been changed to $\mathbf{z}_{1}^{\text{np}}$ encoded by $\mathcal{E}^{\text{np}}$, and the HQ target $\mathbf{z}_{0}$ is compressed from $\mathbf{x}_{0}$ with EP-VAE encoder $\mathcal{E}$.
Furthermore, different from previous bridge models designed for a specific SE task that pursue target waveform reconstruction accuracy, GSR systems are required to generate speech samples with high perceptual quality (Manjunath, 2009; Manocha et al., 2022; Babaev et al., 2024). This key feature motivates us to integrate the human perceptual quality into the entire generation process of VoiceBridge, further enhancing LBM-based GSR quality. Specifically, we perform a post-training stage for both the LBM and the EP-VAE decoder, mitigating their cascading mismatches in inference as well as aligning both the bridge sampling and VAE decoding with human perceptual quality. The post-training objectives for LBMs and the decoder include:
$\mathcal{L}_{\text{pa-gen}}=\mathcal{L}_{\text{bridge}}(\varphi)+\mathcal{L}_{\text{data}}(\mathcal{D}_{\theta}(\hat{\mathbf{z}}_{0,\varphi}(\mathbf{z}_{t},t,\mathbf{z}^{\text{np}}_{1})),\mathbf{x}_{0})+\mathcal{L}_{\text{hf}}(\mathcal{D}_{\theta}(\hat{\mathbf{z}}_{0,\varphi}(\mathbf{z}_{t},t,\mathbf{z}^{\text{np}}_{1})),\mathbf{x}_{0}),$ (7)
where $\mathcal{L}_{\text{hf}}$ denotes a PESQ loss and a UTMOS loss, jointly align the bridge sampling and the decoding processes well with subjective quality judgments. Intuitively, such a post-training procedure helps the bridge model to generate the speech samples on the sub-manifold of the latent space corresponding to the waveform with high perceptual quality, as well as encourages the decoder to generate HQ audio. As the encoder is frozen in this training section, this technique does not affect the good latent properties established by EP-VAE and prior convergence, allowing the utilization of each proposed innovation in VoiceBridge.
@@ 4.1 Experimental Setup
Table 3: Evaluation results on three denoising benchmarks. Best is bolded and second best is underlined. VoiceBridge achieves strong performance on the speech enhancement task compared with both other GSR models and specific SE models.
VB-Demand (Valentini-Botinhao, 2017) Model PESQ ($\uparrow$) SIG ($\uparrow$) BAK ($\uparrow$) OVRL ($\uparrow$) UTMOS ($\uparrow$) WV-MOS ($\uparrow$) NISQA ($\uparrow$) Input 1.972 3.343 3.124 2.694 3.063 2.985 2.998 SBSE 2.140 2.571 3.658 2.946 3.407 3.819 3.976 VoiceFixer 2.454 3.454 4.047 3.174 3.645 4.129 4.433 Resemble-Enhance 2.352 3.454 3.986 3.141 3.678 4.270 4.503 UniverSE++ 3.020 3.486 4.042 3.198 3.968 4.385 4.503 VoiceBridge 2.819 3.471 4.055 3.199 4.288 4.387 4.513 WSJ0-CHiME3 (Barker et al., 2015) Model PESQ ($\uparrow$) SIG ($\uparrow$) BAK ($\uparrow$) OVRL ($\uparrow$) UTMOS ($\uparrow$) WV-MOS ($\uparrow$) NISQA ($\uparrow$) Input 1.248 2.486 1.868 1.795 1.823 0.128 1.520 SGMSE+ 1.723 3.517 3.946 3.191 3.282 3.547 4.294 StoRM 1.826 3.452 4.115 3.214 3.468 3.618 4.611 VoiceFixer 1.493 3.293 4.059 3.052 2.863 3.255 3.921 Resemble-Enhance 1.171 3.524 4.113 3.263 2.749 3.474 4.610 UniverSE++ 1.320 3.173 3.997 2.914 2.984 3.248 3.918 VoiceBridge 1.727 3.536 4.093 3.280 4.440 4.356 4.632 VCTK (Veaux & Yamagishi, 2017) + DEMAND (Thiemann et al., 2013) with SNR=-15dB Model PESQ ($\uparrow$) SIG ($\uparrow$) BAK ($\uparrow$) OVRL ($\uparrow$) UTMOS ($\uparrow$) WV-MOS ($\uparrow$) NISQA ($\uparrow$) Input 1.069 2.091 1.690 1.453 1.352 -1.689 1.129 Voicefixer 1.225 2.558 3.280 2.313 2.423 2.278 2.848 Resemble-Enhance 1.168 3.194 3.396 2.640 3.394 2.440 3.130 UniverSE++ 1.211 2.780 3.202 2.374 2.950 2.772 2.271 VoiceBridge 1.352 3.039 3.690 2.698 3.479 2.755 3.270
We train and evaluate VoiceBridge on GSR task and test its zero-shot performance to downstream tasks, especially out-of-domain (OOD) tasks. Our full training corpus contains approximately 1138 hours of clean speech audio at 48 kHz, constructed by combining multiple public datasets. Distortions are synthesized using publicly available noise datasets and room impulse responses, following previous works (Liu et al., 2022; Zhang et al., 2025b). The compression network is a 156M-parameter Oobleck VAE (Evans et al., 2024b). The LBM is built with a 544M-parameter Transformer backbone (Evans et al., 2024a). The training pipeline of VoiceBridge consists of four stages: (1) pre-training an EP-VAE on clean speech; (2) fine-tuning the EP-VAE encoder to learn a joint neural prior on distorted inputs; (3) training the transformer model to solve the SB in latent space; and (4) jointly fine-tuning the LBM and the decoder to enhance perceptual alignment. We provided detailed information on the VAE compression network, datasets, inference process, and evaluation methods in Appendix DEH, and F, respectively.
@@ 4.2 General Speech Restoration
To comprehensively assess our GSR performance at scale, we first evaluate it on three benchmark datasets following previous works: VoiceFixer-GSR (Liu et al., 2022), DNS-with-Reverb (Reddy et al., 2021a), and DNS-Real-Data (Reddy et al., 2021a). Notably, data from the DNS-Challenge (Reddy et al., 2021a) are excluded in our training data, providing a mismatched training-testing condition for VoiceBridge. We compare VoiceBridge against other GSR models including VoiceFixer (Liu et al., 2022), Resemble-Enhance (Niu, 2024), UniverSE++ (Scheibler et al., 2024) and AnyEnhance (Zhang et al., 2025b).
@@
The GSR evaluation results are shown in Table 1. VoiceBridge achieves the best or the second-best results on almost every metric across all datasets, demonstrating strong multi-degradation restoration capability on both simulated and real-world audios. We further conduct a user study with the publicly available VoiceFixer-GSR and DNS-Real to investigate human preference for each GSR system. The result, as shown in Table 2, demonstrates VoiceBridge’s superior performance, achieving significantly higher Mean Opinion Score (MOS) on both simulated and real audio samples. A case study is shown in Figure 3, where VoiceBridge faithfully generates different frequency components of the ground truth in comparison with other methods. All demonstrated results of VoiceBridge are achieved with 4 latent inference steps, ensuring computational efficiency. We provide more details of inference efficiency in Appendix H.
4.3 Additional Results
Table 4: Codec artifact removal on Encodec-reconstructed audio from the VCTK testset. Though never trained on this task, VoiceBridge exhibits strong zero-shot performance in enhancing speech quality.
VCTK (Veaux & Yamagishi, 2017) reconstructed by Encodec (Défossez et al., 2022) Method PESQ($\uparrow$) WVMOS($\uparrow$) UTMOS($\uparrow$) DNSMOS($\uparrow$) NISQA($\uparrow$) Input 2.210 4.036 2.889 2.822 3.661 Voicefixer 1.984 3.956 3.229 3.003 4.436 Resemble-Enhance 2.059 4.054 3.294 3.063 4.665 UniverSE++ 2.320 4.002 2.954 2.868 3.861 VoiceBridge 2.341 4.109 3.675 3.109 4.452
As VoiceBridge is built with a scalable architecture and aims to achieve high-fidelity GSR at scale, we further evaluate its zero-shot performance on more tasks, and especially, the tasks with out-of-domain (OOD) degradation levels and types. We tested VoiceBridge on the task of Speech Enhancement (SE), comparing it to task-specific models including SGMSE+ (Richter et al., 2023; Welker et al., 2022), StoRM (Lemercier et al., 2023b) and SBSE (Jukić et al., 2024) on benchmarks involving VoiceBank-DEMAND (Valentini-Botinhao, 2017) and WSJ-CHiME (Barker et al., 2015). We additionally curate a VCTK (Veaux & Yamagishi, 2017) + DEMAND (Thiemann et al., 2013) dataset, with a setting at $15\,\mathrm{dB}$ SNR, which is well beyond our training settings, to assess the robustness of VoiceBridge under extreme noise conditions unseen in training. Apart from SE, evaluation for other tasks has also been conducted for bandwidth extension and dereverberation, with details shown in Appendix F. Moreover, as a probabilistic generative model, VoiceBridge captures target distributional information rather than establishing a point-to-point mapping function. Therefore, it shows strong zero-shot performance on OOD restoration tasks that are unseen during the training stage. We evaluate VoiceBridge and other GSR models on codec artifact removal, where all models are tested to repair reconstructed audio from VCTK by the Encodec (Défossez et al., 2022) model operating at 3 kbps bitrate.
The evaluation results of denoising are shown in Table 3, where VoiceBridge outperforms data-space SB models SBSE (Jukić et al., 2024), diffusion models, and flow-matching models on various metrics. Note that the WSJ0 dataset, which the SGMSE+ and StoRM models are trained on, is excluded from our training dataset, exceptionally demonstrating the superiority of VoiceBridge. For the codec artifact removal results, as shown in Table 4, VoiceBridge consistently outperforms baseline models on various metrics, exhibiting the promising zero-shot performance of VoiceBridge on this OOD task.
4.4 Quality Improvement for Speech Generation
@@
Beyond codec artifact removal, we test another application scenario of GSR models, refining the generation results of recent speech and podcast generation methods, namely, further enhancing the perceptual quality of text-to-speech (TTS) models. For validation, we take two recent methods, MaskGCT (Wang et al., 2024b) and MoonCast (Ju et al., 2025), as the TTS baseline methods to synthesize speech on the Seed-TTS (Boson AI, 2025; Peng et al., 2025) benchmark, and then apply VoiceBridge and other GSR baselines to refine the generated speech samples. We evaluate using both GSR and TTS metrics. As shown in Table 5, VoiceBridge reliably improves the quality of recent speech generation methods in a zero-shot manner. Importantly, other GSR models may fail to achieve high-fidelity GSR at scale, and therefore they are not able to reliably refine recent TTS methods, while VoiceBridge consistently refines these methods.
Table 5: Enhancement results on refining the generated speech by MoonCast and MaskGCT on the Seed-TTS benchmark. VoiceBridge exhibits consistent improvement on generated speech quality.
Seed-TTS-Eval (Anastassiou et al., 2024) with MoonCast (Ju et al., 2025) Method WVMOS($\uparrow$) UTMOS($\uparrow$) SIG($\uparrow$) BAK($\uparrow$) OVRL($\uparrow$) NISQA($\uparrow$) WER($\downarrow$) Input 3.971 3.623 3.451 3.880 3.120 3.922 5.93% Voicefixer 3.832 3.481 3.433 3.993 3.145 4.004 5.51% Resemble-Enhance 3.612 3.291 3.424 3.972 3.176 4.121 12.88% UniverSE++ 3.747 3.446 3.459 3.957 3.143 4.217 8.59% VoiceBridge 4.077 3.807 3.482 4.019 3.208 4.443 5.40% Seed-TTS-Eval (Anastassiou et al., 2024) with MaskGCT (Wang et al., 2024b) Method WVMOS($\uparrow$) UTMOS($\uparrow$) SIG($\uparrow$) BAK($\uparrow$) OVRL($\uparrow$) NISQA($\uparrow$) WER($\downarrow$) Input 3.673 3.427 3.529 3.793 3.122 3.964 5.63% Voicefixer 3.483 3.288 3.478 4.012 3.186 4.349 6.02% Resemble-Enhance 3.818 3.594 3.580 4.071 3.312 4.691 7.78% UniverSE++ 3.324 3.277 3.475 4.016 3.180 4.308 7.21% VoiceBridge 4.156 4.040 3.565 4.088 3.311 4.764 5.69%
@@ 4.5 Ablation Studies
We conduct an ablation study on our designs by training four model variants: with or without EP, and with or without the joint neural prior. All variants share the same latent bridge and decoder training pipeline for fair comparison. As shown in Figure 4(d), the baseline model without EP and joint neural prior consistently underperforms the other variants. Introducing the joint neural prior or adding the EP constraint leads to a noticeable improvement. The model with both techniques achieves the best overall performance, demonstrating the complementary benefits of these two mechanisms.
@@
Further, we conduct another ablation study on the training objective and strategy during the perceptual aware fine-tuning stage. We fine-tune the bridge model in 6 settings with different combinations of additional data-space reconstruction loss (R.), human feedback loss (H.), adversarial and feature matching loss (AF.). Additionally, an extra experiment is done with frozen decoder, tuning the bridge only (B.). As shown in Table 6, combining these loss terms to jointly fine-tune both the bridge and decoder is exceptionally important for achieving a LBM-based GSR system with high perceptual quality. More details are shown in Appendix I.
Table 6: Ablation study on the perceptual aware finetuning stage. Cont. denotes continue training with the original bridge loss. R, H, A, F denotes Reconstruction, Human-feedback, Adversarial, and Feature-matching loss on the reconstructed data space. B denotes finetuning latent bridge model only with the decoder frozen.
Cont. R RH RAF RHAF RHAF, B WV-MOS 4.153 3.972 4.252 3.709 4.366 4.158 NISQA 4.067 2.301 3.167 3.091 4.172 3.487
As shown in Table 15, conducting post-training with the decoder frozen makes little improvement to the restoration system. Adding only the reconstruction loss or with the GAN loss for the joint finetuning also have negligible or even negative effects. Interestingly, adding the human feedback term $\mathcal{L}_{\text{hf}}$ has a significant impact on the corresponding metrics, PESQ and UTMOS, but causes a decrease in other metrics. Adding all objectives achieves a consistent and signficant improvement in all metrics, demonstrating a real advancement in perceptual quality. More empirical discoveries and detailed analysis can be seen in Appendix I

### v3

@@ 3.4 From Denoiser to Generator
After training the EP-VAE encoder $\mathcal{E}$ with Equation 5 to encode HQ target, its fine-tuned version $\mathcal{E}^{\text{np}}$ with Equation 6 for diverse LQ priors, and the shared decoder $\mathcal{D}$, we can already create a structured latent space where the different LQ priors have converged. Then, we train the bridge transformer model that maps the joint neural prior and the HQ target using Equation 3, while the LQ prior and condition $\mathbf{z}_{1}$ have been changed to $\mathbf{z}_{1}^{\text{np}}$ encoded by $\mathcal{E}^{\text{np}}$, and the HQ target $\mathbf{z}_{0}$ is compressed from $\mathbf{x}_{0}$ with EP-VAE encoder $\mathcal{E}$.
Although the latent bridge model (LBM) and the VAE decoder are trained sequentially, their objectives are not perfectly aligned: the LBM is optimized to predict target latents under the bridge loss, while the decoder is optimized to reconstruct waveforms from encoder latents. At inference, this discrepancy can lead to cascading mismatch: small deviations in LBM-predicted latents may be amplified by the decoder, degrading perceptual quality. To mitigate this issue, we introduce a post-training stage that jointly fine-tunes the LBM and the decoder, while keeping the encoder fixed. This preserves the latent geometry learned by the encoder, and simultaneously calibrates the decoder to the LBM-induced latent distribution.
To ensure fine supervision from the data domain, we introduce the data-level reconstruction loss $\mathcal{L}_{\text{data}}$ in VAE training to the objectives of the post-training process. Meanwhile, as GSR systems are required to generate speech samples with high perceptual quality (Manjunath, 2009; Manocha et al., 2022; Babaev et al., 2024), we incorporate a perceptual metric aligned loss $\mathcal{L}_{\text{perc}}$ utilizing the PESQ (Rix et al., 2001) and UTMOS (Saeki et al., 2022) metrics. However, including only this term causes severe overfitting problems (de Oliveira et al., ), as the decoder falls into a local minima of these metrics and forget faithful reconstruction. To tackle this issue, we activate the discriminator $\omega$ used in VAE training, and incorporate an adversarial objective $\mathcal{L}_{\text{GAN}}$ to the joint post-training process. The perceptual objective aligns the generation process with human perceptual quality, and the discriminator is reponsible to spot artifacts caused by optimizing the perceptual indicators and prevent overfitting.
Additionally, the adversarial objective surprisingly can transform the LBM from a multi-step denoiser into a single-step generator, enabling one-step inference without distillation. In the LBM training stage guided by the MSE loss 3, the model learns an optimal $\hat{\mathbf{z}}_{0,\varphi}$ estimating
$\arg\min_{\hat{\mathbf{z}}_{0,\varphi}}\ \mathbb{E}\left[\|\mathbf{z}_{0}-\hat{\mathbf{z}}_{0,\varphi}(\mathbf{z}_{t},\mathbf{z}_{1}^{\text{np}})\|_{2}^{2}\right]=\mathbb{E}[\mathbf{z}_{0}\mid\mathbf{z}_{t},\mathbf{z}_{1}^{\text{np}}],$ (7)
The one-step prediction of $\hat{\mathbf{z}}_{0,\varphi}$ is not a sample from the target distribution, but a conditional expectation symbolizing the denoising target. However, the adversarial loss aims to $\min_{\varphi}\ \max_{\omega}\ \mathcal{L}_{\text{GAN}}(\varphi,\omega)$, leading $\hat{\mathbf{z}}_{0,\varphi}$ from the conditional expectation toward the conditional distribution
$p_{\varphi}(\mathbf{z}_{0}\mid\mathbf{z}_{t},\mathbf{z}_{1}^{\text{np}})=p_{\text{data}}(\mathbf{z}_{0}\mid\mathbf{z}_{t},\mathbf{z}_{1}^{\text{np}}),$ (8)
at the global optimum, which gives the model one-step restoration ability. Detailed derivations are in Appendix E
Technically, we perform a post-training stage for both the LBM and the EP-VAE decoder, mitigating their cascading mismatches in inference as well as aligning both the bridge sampling and VAE decoding with human perceptual quality. The post-training objectives for LBMs and the decoder include:
$\mathcal{L}_{\text{pt}}=$ $\mathcal{L}_{\text{bridge}}(\varphi)+\mathcal{L}_{\text{data}}(\theta,\varphi)$
$+$ $\mathcal{L}_{\text{perc}}(\theta,\varphi)+\mathcal{L}_{\text{GAN}}(\theta,\varphi,\omega),$ (9)
where $\mathcal{L}_{\text{perc}}$ denotes a PESQ loss and a UTMOS loss, and $\mathcal{L}_{\text{data}}$ is the data level MSE loss and $\mathcal{L}_{\text{GAN}}$ the adversarial loss. These three losses are all calculated between decoded signal $\hat{\mathbf{z}}_{0,\varphi}(\mathbf{z}_{t},t,\mathbf{z}^{\text{np}}_{1})$ and ground truth $\mathbf{x}_{0}$, with weights updated on LBM $\varphi$, decoder $\theta$, and discriminator $\omega$ differently.
3.5 Training Pipeline Overview
In actual training, VoiceBridge integrates EP-VAE and the joint neural prior, and achieves denoiser-to-generator alignment through a four-stage training pipeline: (1) pre-training an EP-VAE on clean speech with Equation 4; (2) fine-tuning the EP-VAE encoder to learn a joint neural prior on distorted inputs with Equation 6; (3) training the transformer model to solve the SB in latent space with Equation 3 and (4) jointly fine-tuning the LBM and the decoder to enhance perceptual alignment with Equation 3.4.
Table 2: Enhancement results on OOD tasks, including codec artifact removal and TTS quality improvement. VoiceBridge exhibits consistent improvement on generated speech quality.
Task Dataset Data Domain Model($\uparrow$) WVMOS($\uparrow$) UTMOS($\uparrow$) DNSMOS($\uparrow$) NISQA($\uparrow$) PESQ($\uparrow$) WER($\downarrow$) Codec Artifact Removal VCTK (Encodec) In-Domain Data Input 4.036 2.889 2.822 3.661 2.210 - VoiceFixer 3.956 3.229 3.003 4.436 1.984 - Resemble-Enhance 4.054 3.294 3.063 4.665 2.059 - UniverSE++ 4.002 2.954 2.868 3.861 2.320 - VoiceBridge 4.131 3.683 3.132 4.469 2.361 - TTS Quality Improvement Seed-TTS (MoonCast) Out-Domain Data Input 3.971 3.623 3.120 3.922 - 5.93% VoiceFixer 3.832 3.481 3.145 4.004 - 5.51% Resemble-Enhance 3.612 3.291 3.176 4.121 - 12.88% UniverSE++ 3.747 3.446 3.143 4.217 - 8.59% VoiceBridge 4.098 3.821 3.227 4.464 - 5.40% Seed-TTS (MaskGCT) Out-Domain Data Input 3.673 3.427 3.122 3.964 - 5.63% VoiceFixer 3.483 3.288 3.186 4.349 - 6.02% Resemble-Enhance 3.818 3.594 3.312 4.691 - 7.78% UniverSE++ 3.324 3.277 3.180 4.308 - 7.21% VoiceBridge 4.172 4.051 3.326 4.769 - 5.66%
@@ 4.1 Experimental Setup
We train and evaluate VoiceBridge on GSR task, and also test its zero-shot generalization to out-of-domain (OOD) data and tasks. Our full training corpus contains approximately 1138 hours of clean speech audio at 48 kHz, constructed only by combining public datasets. Distortions are synthesized using publicly available noise datasets and room impulse responses, following previous works (Liu et al., 2022; Zhang et al., 2025b). The compression network adopts the Oobleck (Evans et al., 2024b) arthitecture. The LBM is built with a 544M-parameter Transformer backbone (Evans et al., 2024a). The training process is described in Section 3.5. Only one inference step is used. We provide detailed information on the VAE compression network, datasets, and evaluation methods in Appendix DF, and G, respectively.
@@ 4.2 General Speech Restoration
We first evaluate VoiceBridge’s performance on in-domain restoration tasks, including simulated GSR task and downstream single-task speech enhancement. Following previous works, we use VoiceFixer-GSR (Liu et al., 2022), DNS-with-Reverb (Reddy et al., 2021a) as the simulated GSR benchmarks. Notably, data from the DNS-Challenge (Reddy et al., 2021a) are excluded in our training data, providing a mismatch in training test conditions for VoiceBridge. We compare VoiceBridge against other GSR models including Miipher (Koizumi et al., 2023), VoiceFixer (Liu et al., 2022), Resemble-Enhance (Niu, 2024), UniverSE++ (Scheibler et al., 2024) and AnyEnhance (Zhang et al., 2025b). Despite these models, we further compare VoiceBridge to close-sourced models with there reported results or using there demo samples in Appendix G.
@@
The GSR evaluation results are shown in Table 1. VoiceBridge achieves the best or the second-best results on almost every metric across all datasets, demonstrating strong multi-degradation restoration capability on both simulated and real-world audios. We further conduct a user study with VoiceFixer-GSR and DNS-Real to investigate human preference for each GSR system. The result, as shown in Table 3, demonstrates VoiceBridge’s superior performance, achieving significantly higher Mean Opinion Score (MOS) on both simulated and real audio samples. A case study is shown in Figure 3, where VoiceBridge faithfully generates different components of the ground truth with four steps in comparison with other methods.
Further, as a GSR model, VoiceBridge naturally handles all kinds of restoration tasks within the data distribution. We test VoiceBridge on the task of traditional Speech Enhancement (SE), comparing it to task-specific models including SGMSE+ (Richter et al., 2023; Welker et al., 2022), StoRM (Lemercier et al., 2023b) and SBSE (Jukić et al., 2024) on benchmarks involving VoiceBank-DEMAND (Valentini-Botinhao, 2017) and WSJ-CHiME (Barker et al., 2015). The results are shown in Table 1, where VoiceBridge outperforms data-space SB models SBSE (Jukić et al., 2024), diffusion models, and flow-matching models on various metrics. Note that the WSJ0 dataset, which the SGMSE+ and StoRM models are trained on, is excluded from our training dataset, exceptionally demonstrating the superiority of VoiceBridge. Apart from SE, evaluation for other tasks has also been conducted, with details in Appendix G.
4.3 Out of Domain Tasks
@@
Moreover, as a probabilistic generative model, VoiceBridge captures target distributional information rather than establishing a point-to-point mapping function. Therefore, it shows strong zero-shot ability on OOD restoration tasks that are unseen during the training stage. We evaluate VoiceBridge and other GSR models on codec artifact removal, where all models are tested to repair reconstructed audio from VCTK by the Encodec (Défossez et al., 2022) model operating at 3 kbps bitrate. Beyond that, We originally propose another application scenario of GSR models: further refining the perceptual quality of generation results of text-to-speech (TTS) models. We take two recent methods, MaskGCT (Wang et al., 2024b) and MoonCast (Ju et al., 2025), as the TTS baseline methods to synthesize speech on the Seed-TTS (Boson AI, 2025; Peng et al., 2025) benchmark, and then apply VoiceBridge and other GSR baselines to refine the generated speech samples. We evaluate using both GSR and TTS metrics.
As Shown in Table 2, VoiceBridge consistently outperforms baseline models on various metrics, exhibiting the zero-shot performance of VoiceBridge on OOD tasks. Importantly, in Table 2, other GSR models may fail to further enhance the model-generated speech, while VoiceBridge reliably improves the quality. This demonstrates VoiceBridge’s strong capability of comprehensively refining audio quality.
@@ 4.4 Ablation Studies
We conduct an ablation study on our designs by training four model variants: with or without EP, and with or without the joint neural prior. All variants share the same latent bridge and decoder training pipeline for fair comparison. As shown in Figure 4(d), introducing the joint neural prior or adding the EP constraint leads to a noticeable improvement, while adding both achieves the best overall performance, demonstrating the complementary benefits of these two mechanisms.
@@
Table 4: Evaluation results on VoiceFixer-GSR (Liu et al., 2022) for different post-training processes of VoiceBridge.
Voicefixer-GSR (Liu et al., 2022) Finetuned Module Finetuning Objective PESQ ($\uparrow$) DNSMOS ($\uparrow$) UTMOS ($\uparrow$) WV-MOS ($\uparrow$) NISQA ($\uparrow$) Bridge $\mathcal{L}_{\text{lat.}}$ 2.067 3.031 3.645 4.153 4.067 $\mathcal{L}_{\text{lat.}}+\mathcal{L}_{\text{rec}}$ 2.125 3.049 3.649 4.207 3.524 $\mathcal{L}_{\text{lat.}}+\mathcal{L}_{\text{rec}}+\mathcal{L}_{\text{gan}}$ 2.058 3.069 3.533 4.055 3.362 $\mathcal{L}_{\text{lat.}}+\mathcal{L}_{\text{rec}}+\mathcal{L}_{\text{perc}}$ 2.171 3.038 3.625 4.122 3.466 $\mathcal{L}_{\text{lat.}}+\mathcal{L}_{\text{rec}}+\mathcal{L}_{\text{gan}}+\mathcal{L}_{\text{perc}}$ 2.182 3.049 3.599 4.158 3.487 Bridge ++ Decoder $\mathcal{L}_{\text{lat.}}+\mathcal{L}_{\text{rec}}$ 2.297 2.952 3.588 3.972 2.301 $\mathcal{L}_{\text{lat.}}+\mathcal{L}_{\text{rec}}+\mathcal{L}_{\text{gan}}$ 2.169 2.938 3.475 3.709 3.091 $\mathcal{L}_{\text{lat.}}+\mathcal{L}_{\text{rec}}+\mathcal{L}_{\text{perc}}$ 3.014 3.025 4.267 4.252 3.167 $\mathcal{L}_{\text{lat.}}+\mathcal{L}_{\text{rec}}+\mathcal{L}_{\text{gan}}+\mathcal{L}_{\text{perc}}$ 2.722 3.091 4.228 4.366 4.172
Further, we ablate the training objectives and update strategy in the post-training stage. Starting from the pre-trained LBM, we run 9 post-training variants that differ in (i) the loss objective and (ii) whether the decoder is updated. We consider two groups: (1) decoder frozen, where we fine-tune only the bridge using a data-space loss between decoded audio and clean audio; and (2) joint post-training, where the bridge and decoder are fine-tuned together.
We evaluate the loss combinations in Table 4. As a reference baseline, we continue training with the original latent-space MSE objective $\mathcal{L}_{\text{lat.}}$ (Line 1). For post-training in the data domain, we use MRSTFT reconstruction loss $\mathcal{L}_{\text{rec}}$ (the same loss used in VAE training) as the base objective (Lines 2 and 6). We then add the GAN loss $\mathcal{L}_{\text{gan}}$ (adversarial + feature matching) and the perceptual term $\mathcal{L}_{\text{perc}}$, first individually (Lines 3,4,7,8) and then together (Lines 5 and 9). Each objective variant is tested in both update strategies (decoder frozen vs. joint). All hyperparameters follow the VAE training setup in Appendix D. All results use 4 inference steps.
As shown in Table 4, freezing the decoder yields little improvement regardless of the objective. Under joint post-training, using $\mathcal{L}_{\text{rec}}$ alone, or $\mathcal{L}_{\text{rec}}$ with $\mathcal{L}_{\text{gan}}$, produces negligible or even negative gains. This may leads to the change in the MSE objective, causing approximation error in the bridge transport trajectory. In contrast, adding only $\mathcal{L}_{\text{perc}}$ substantially increases PESQ and UTMOS but degrades other metrics, consistent with metric-overfitting behavior reported in PESQetarian (de Oliveira et al., ). The generated audio shares a common audible artifact, degrading listening experience. Intuitively, since the evaluator is fixed, the decoder can learn “hacks” that exploit the perceptual scorer rather than improving true audio quality.
Combining $\mathcal{L}_{\text{gan}}$ and $\mathcal{L}_{\text{perc}}$ prevents such metric hacking. The final setting (Line 9) yields a consistent and significant improvement across all metrics, indicating a genuine gain in perceptual quality. A plausible explanation is that overfitting-induced artifacts are readily detected by the discriminator, forcing the model to raise perceptual scores by improving generation quality.
We also conduct further ablation experiments regarding the design space of bridge modeling and EP-VAE training objectives. Results are shown in Appendix I due to page limit.